# Asynchronous Predictive Counterfactual Regret Minimization+ Algorithm in Solving Extensive-Form Games

### v1

@@ Abstract
Counterfactual Regret Minimization (CFR) algorithms are widely used to compute a Nash equilibrium (NE) in two-player zero-sum imperfect-information extensive-form games (IIGs). Among them, Predictive CFR+ (PCFR+) is particularly powerful, achieving an exceptionally fast empirical convergence rate via the prediction in many games. However, the empirical convergence rate of PCFR+ would significantly degrade if the prediction is inaccurate, leading to unstable performance on certain IIGs. To enhance the robustness of PCFR+, we propose a novel variant, Asynchronous PCFR+ (APCFR+), which employs an adaptive asynchronization of step-sizes between the updates of implicit and explicit accumulated counterfactual regrets to mitigate the impact of the prediction inaccuracy on convergence. We present a theoretical analysis demonstrating why APCFR+ can enhance the robustness. Finally, we propose a simplified version of APCFR+ called Simple APCFR+ (SAPCFR+), which uses a fixed asynchronization of step-sizes to simplify the implementation that only needs a single-line modification of the original PCFR+. Interestingly, SAPCFR+ achieves a constant-factor lower theoretical regret bound than PCFR+ in the worst case. Experimental results demonstrate that (i) both APCFR+ and SAPCFR+ outperform PCFR+ in most of the tested games, as well as (ii) SAPCFR+ achieves a comparable empirical convergence rate with APCFR+.
Machine Learning, ICML
@@ 1 Introduction
Imperfect-information extensive-form games (IIGs) are foundational models to capture interactions among multiple agents in sequential settings with hidden information. IIGs are widely used to simulate real-world scenarios such as medical treatment (Sandholm, 2015), security games (Lisỳ et al., 2016), cybersecurity (Chen et al., 2017), and recreational games (Brown & Sandholm, 2018, 2019b). To address IIGs, a primary goal is to compute a Nash equilibrium (NE), where no player can unilaterally improve its payoff by deviating from the equilibrium.
As with much of the literature on solving IIGs, we focus on computing an NE in two-player zero-sum IIGs. The most widely used method for computing an NE in these IIGs is Counterfactual Regret Minimization (CFR)  (Zinkevich et al., 2007; Lanctot et al., 2009; Tammelin, 2014; Brown & Sandholm, 2019a; Farina et al., 2021, 2019; Liu et al., 2021, 2023; Meng et al., 2023; Farina et al., 2023; Xu et al., 2022, 2024b, 2024a; Zhang et al., 2024), as evidenced by their success in superhuman game AIs (Bowling et al., 2015; Moravčík et al., 2017; Brown & Sandholm, 2018, 2019b; Pérolat et al., 2022). The key insight of CFR algorithms is to decompose the total regret over the game into a sum of counterfactual regrets associated within information sets (infosets) and employ a local regret minimizer to minimize counterfactual regrets within each infoset.
Many technologies have been proposed to improve the empirical convergence rate of CFR algorithms. For example, Counterfactual Regret Minimization+ (CFR+(Tammelin, 2014) replaces the local regret minimizer—Regret Matching (RM) (Hart & Mas-Colell, 2000; Gordon, 2006)—used in vanilla CFR (Zinkevich et al., 2007) with Regret Matching+ (RM+). CFR+ improves the empirical convergence rate by ensuring that the accumulated counterfactual regrets remain non-negative. Subsequently, Farina et al. (2021) introduce Predictive CFR+ (PCFR+), an improved variant of CFR+.
PCFR+ significantly outperforms other CFR algorithms in many IIGs by using the prediction. Specifically, PCFR+ maintains two types of accumulated counterfactual regrets: the implicit and the explicit. As shown in Figure 1, at each iteration $t$, PCFR+ uses the prediction and the observed instantaneous counterfactual regret $\bm{r}^{t}_{I}$ to derive the new explicit accumulated counterfactual regret $\hat{\bm{R}}^{t}_{I}$ and the new implicit counterfactual regret $\bm{R}^{t+1}_{I}$, respectively. If the prediction aligns with the observed instantaneous counterfactual regret $\bm{r}^{t}_{I}$, the theoretical convergence rate of PCFR+ can be improved from $O(1/\sqrt{T})$ of CFR+ to $O(1/T)$ (Farina et al., 2021). However, PCFR+ sets the instantaneous counterfactual regret $\bm{r}^{t-1}_{I}$ observed at iteration $t-1$ as the prediction at iteration $t$. This operation may cause inaccurate prediction on certain IIGs, which harms the empirical convergence rate of PCFR+. As noted by Farina et al. (2021), PCFR+ performs worse than other CFR algorithms in Leduc Poker, yet significantly surpasses them in Battleship (3). This aligns with the results in Figure 2: the gap between predicted and observed instantaneous counterfactual regret decreases slowly in Leduc Poker but diminishes rapidly in Battleship (3), validating our hypothesis.
@@
Figure 1: Comparison between the update rules of PCFR+ and APCFR+, with differences highlighted in red.
@@
Figure 2: Dynamics of accumulated inaccuracy in PCFR+ between predicted and observed instantaneous counterfactual regrets in Leduc Poker and Battleship (3). This accumulated inaccuracy is related to theoretical convergence rate of PCFR+ (Theorem 4.1). The values on the y-axis are normalized to the range [0, 1].
@@
To enhance the robustness of PCFR+, we propose a novel variant of PCFR+, termed Asynchronous PCFR+ (APCFR+). Similar to PCFR+, APCFR+ leverages the prediction to improve the convergence rate, but it mitigates the impact of the prediction inaccuracy on convergence. Specifically, as illustrated in Figure 1, APCFR+ utilizes an adaptive asynchronization mechanism for step-sizes between implicit and explicit accumulated counterfactual regret updates, which dynamically reduces the step-size when updating via the prediction. We prove that when the step-size for updating the explicit accumulated counterfactual regret via the prediction at iteration $t$ is set to ${1}/{(1+\alpha^{t})}$ for APCFR+, where $\alpha^{t}\geq 0$, the effect of the prediction inaccuracy on the convergence rate for APCFR+ is reduced by a factor of $1+\alpha^{t}$ compared to PCFR+. Therefore, APCFR+ mitigates the impact of the prediction inaccuracy on the convergence rate. Then, through the theoretical analysis of APCFR+, we propose an automatic learning mechanism for $\alpha^{t}$, eliminating the need for fine-tuning parameters across different games. To the best of our knowledge, we are the first to use different step-sizes for updating implicit and explicit accumulated counterfactual regrets.
We also introduce a simplified version of APCFR+, called Simple APCFR+ (SAPCFR+), which sets $\alpha^{t}$ as a constant to simplify the implementation that SAPCFR+ can be implemented with a single-line modification to the PCFR+ code. Notably, the constant is derived from the theoretical analysis of APCFR+, guaranteeing that SAPCFR+ achieves a constant-factor improvement in the worst-case theoretical convergence rate compared to PCFR+.
@@
We conduct extensive experimental evaluations of APCFR+ across nine instances from five standard IIG benchmarks, including Kuhn Poker, Leduc Poker, Goofspiel, Liar’s Dice, and Battleship, as well as two heads-up no-limit Texas Hold’em (HUNL) Subgames generated by the top poker agent, Libratus (Brown & Sandholm, 2018). The experiments demonstrate that our algorithms outperforms PCFR+ in almost all tested games, and achieves an empirical convergence rate comparable to that of PCFR+ at a minimum. Moreover, we observe that SAPCFR+ achieves comparable empirical convergence rate with APCFR+.
@@ 2 Related Work
We focus on CFR algorithms (Zinkevich et al., 2007; Lanctot et al., 2009; Tammelin, 2014; Brown & Sandholm, 2019a; Farina et al., 2021, 2019; Liu et al., 2021; Pérolat et al., 2021; Liu et al., 2023; Meng et al., 2023; Farina et al., 2023; Xu et al., 2022, 2024b, 2024a; Zhang et al., 2024), the most widely used method for learning an NE in two-player zero-sum IIGs, as evidenced by their success in superhuman game AIs (Bowling et al., 2015; Moravčík et al., 2017; Brown & Sandholm, 2018, 2019b; Pérolat et al., 2022).
The key insight of CFR algorithms is the decomposition of the regret over the game into the sum of counterfactual regrets associated within infosets. The vanilla CFR algorithm, introduced by Zinkevich et al. (2007), employs RM (Hart & Mas-Colell, 2000) as the local regret minimizer. To improve the empirical convergence rate of CFR, it is common to design more effective local regret minimizers, as the selection of the local regret minimizers has a significant impact on the overall performance of the CFR algorithm. Examples include RM+ (Bowling et al., 2015), Discounted RM (DRM) (Brown & Sandholm, 2019a), and PRM+ (Farina et al., 2021), which correspond to CFR+ (Bowling et al., 2015), DCFR (Brown & Sandholm, 2019a), and PCFR+ (Farina et al., 2021), respectively. PCFR+ can demonstrate a extremely faster empirical convergence rate than other CFR variants. However, as shown in its original paper, PCFR+ is outperformed by CFR+ and DCFR even on standard IIG benchmarks like Leduc Poker and Liar’s Dice.
To improve the robustness of PCFR+, Farina et al. (2023) propose Stable PCFR+ and Smooth PCFR+. These algorithms improve the robustness by addressing the instability, i.e., rapid strategy fluctuations across iterations, via ensuring the lower bound of the 1-norm of accumulated counterfactual regrets exceeds a positive constant. However, these algorithms never outperform PCFR+ even they achieve a faster theoretical convergence rate than PCFR+. APCFR+ does not focus on addressing the instability, but instead aims to mitigate the impact of the prediction inaccuracy on the convergence to improve the robustness. In our experiments, APCFR+ consistently outperforms Stable PCFR+ and Smooth PCFR+ in all tested games.
@@ 3 Preliminaries
[…]
@@
Behavioral strategy. In this paper, we present the strategy via behavioral strategy. This strategy $\sigma_{i}$ is defined on each infoset. For any infoset $I\in\mathcal{I}_{i}$, the probability for an action $a\in A(I)$ is denoted by $\sigma_{i}(I,a)$. We use $\sigma_{i}(I)=[\sigma_{i}(I,a)|a\in A(I)]\in\Delta^{|A(I)|}$ to denote the strategy at infoset $I$, where $\Delta^{|A(I)|}$ is a $(|A(I)|-1)$-dimension simplex. If every player follows the strategy profile $\sigma=[\sigma_{0};\sigma_{1}]$ and reaches infoset $I$, the reaching probability is denoted by $\pi^{{\sigma}}(I)$. The contribution of $i$ to this probability is $\pi_{i}^{\sigma}(I)$ and $\pi_{-i}^{\sigma}(I)$ for other than $i$, where $-i$ denotes the players other than $i$. In IIGs, $u_{i}(\sigma_{i},\sigma_{-i})=\sum_{z\in\mathcal{Z}}u_{i}(z)\pi^{\sigma}(z)$.
@@
Nash equilibrium (NE). NE denotes a rational behavior where no player can benefit by unilaterally deviating from the equilibrium. For any player, her strategy is the best-response to the strategies of others. Formally, for any NE strategy profile $\sigma^{*}$ and $i\in\mathcal{N}$, it holds that $u_{i}(\sigma^{*}_{i},\sigma^{*}_{-i})\geq u_{i}(\sigma_{i},\sigma^{*}_{-i})$ for all $\sigma$. A widely used metric to measure the distance from the given strategy profile $\bm{x}$ to NE is the exploitability, which is defined as $\epsilon(\sigma)=\sum_{i\in\mathcal{N}}\max_{\sigma^{\prime}_{i}}(u_{i}(\sigma_{i},\sigma_{-i})-u_{i}(\sigma^{\prime}_{i},\sigma_{-i}))/|\mathcal{N}|$. If $\epsilon(\sigma)=\delta$, then $\sigma$ is called as an $\delta$-NE.
@@
Computing an NE via regret minimization algorithms. To compute an NE in IIGs, a common used method is regret minimization algorithms (Rakhlin & Sridharan, 2013a, b; Hazan et al., 2016; Joulani et al., 2017). For any sequence of strategies $\sigma_{i}^{1},\cdots,\sigma_{i}^{T}$ of player $i$, player $i$’s regret is $R^{T}_{i}=\max_{\sigma_{i}}\sum^{T}_{t=1}u_{i}(\sigma_{i},\sigma^{t}_{-i})-\sum^{T}_{t=1}u_{i}(\sigma^{t}_{i},\sigma^{t}_{-i})$ for all sequence $\sigma_{-i}^{1},\cdots,\sigma_{-i}^{T}$. Regret minimization algorithms are algorithms ensuring $R^{T}_{i}$ grows sublinearly. If each player follows a regret minimization algorithm, then their average strategy converges to the set of the NE in two-player zero-sum IIGs. Formally, assume the regret of each player $i$ is $R^{T}_{i}$, then it hods that
$\footnotesize\epsilon(\bar{\sigma})=\epsilon(\bar{\sigma}_{0},\bar{\sigma}_{1})\leq\frac{\sum_{i\in\mathcal{N}}R^{T}_{i}}{|\mathcal{N}|T},$ (1)
where $\bar{\sigma}_{i}(I)=\sum_{t=1}^{T}\pi^{\sigma^{t}}_{i}(I){\sigma^{t}_{i}}(I)/\sum_{t=1}^{T}\pi^{\sigma^{t}}_{i}(I)$ for all $i\in\mathcal{N}$ and $I\in\mathcal{I}_{i}$.
@@
Counterfactual Regret Minimization (CFR) framework. This framework (Zinkevich et al., 2007; Farina et al., 2019; Liu et al., 2021; Farina et al., 2023) is designed to compute an NE of two-player zero-sum IIGs. Instead of directly minimizing the global regret, this framework decomposes the regret to each infoset and independently minimizes the local regret within each infoset. Let $\sigma^{t}$ be the strategy profile at iteration $t$. CFR algorithms compute the counterfactual value at infoset $I$ for action $a$ as
$\footnotesize v^{\sigma^{t}}(Ia)=\sum_{h\in I}\sum_{z\in\mathcal{Z}_{ha}}\pi^{\sigma^{t}}_{-i}(h)\pi^{\sigma^{t}}(ha,z)u_{i}(z),$ (2)
where $\pi^{\sigma^{t}}(ha,z)$ denotes the probability from $ha$ to $z$ if all players play according to $\sigma^{t}$ and $\mathcal{Z}_{ha}$ is the set of the leaf nodes that are reachable after choosing action $a$ at history $h$. For any infoset $I$, the counterfactual regret is
$\footnotesize R^{T}(I)=\max_{a\in A(I)}\sum_{t=1}^{T}v^{\sigma^{t}}(Ia)-\sum_{t=1}^{T}\sum_{a\in A(I)}\sigma^{t}_{i}(Ia)v^{\sigma^{t}}(Ia).$ (3)
Zinkevich et al. (2007) show that the regret over the game $R^{T}_{i}=\max_{\sigma_{i}}\sum^{T}_{t=1}u_{i}(\sigma_{i},\sigma^{t}_{-i})-\sum^{T}_{t=1}u_{i}(\sigma^{t}_{i},\sigma^{t}_{-i})$ is less than the sum of the counterfactual regrets within infosets:
$\footnotesize R^{T}_{i}\leq\sum_{I\in\mathcal{I}_{i}}R^{T}(I).$ (4)
So any regret minimization algorithms can be used as the local regret minimizer to minimize the regret $R^{T}(I)$ over each infoset to minimize the regret $R^{T}_{i}$.
Vanilla Counterfactual Regret Minimization (Vanilla CFR). The first CFR algorithm is proposed by Zinkevich et al. (2007), which uses Regret Matching (RM) (Hart & Mas-Colell, 2000; Gordon, 2006) as the local regret minimizer. Formally, at each iteration $t$, vanilla CFR updates its accumulated counterfactual regret ${\bm{R}}^{t}_{I}$ (notably, it is a vector while $R^{T}_{i}$ is a scalar) at infoset $I$ via
${\bm{R}}^{t+1}_{I}={\bm{R}}^{t}_{I}+\bm{r}^{t}_{I},$ (5)
where $\bm{r}^{t}_{I}=v^{\sigma^{t}}(I)-\langle v^{\sigma^{t}}(I),\sigma^{t}_{i}(I)\rangle\bm{1}$ is the instantaneous counterfactual regret. Then, vanilla CFR gets the new strategy via the regret-matching operator
$\sigma^{t+1}_{i}(I)=\frac{[{\bm{R}}^{t+1}_{I}]^{+}}{\|[{\bm{R}}^{t+1}_{I}]^{+}\|_{1}},$ (6)
where $i=P(I)$ and $[\cdot]^{+}=\max(\cdot,\bm{0})$.
Counterfactual Regret Minimization+ (CFR+). To improve the empirical convergence rate of vanilla CFR, Tammelin (2014) propose a variant of vanilla CFR called CFR+, which utilizes Regret Matching+ (RM+(Tammelin, 2014) as the local regret minimizer. The key insight of RM+ is to ensure that ${\bm{R}}^{t}_{I}\geq\bm{0}$ at each iteration $t$ and infoset $I$. Formally, at each iteration $t$, CFR+ updates its accumulated counterfactual regret ${\bm{R}}^{t}_{I}$ at infoset $I$ according to
${\bm{R}}^{t+1}_{I}=[{\bm{R}}^{t}_{I}+\bm{r}^{t}_{I}]^{+}.$ (7)
Then, as done in CFR, CFR+ gets strategies via the regret-matching operator
$\sigma^{t+1}_{i}(I)=\frac{[{\bm{R}}^{t+1}_{I}]^{+}}{\|[{\bm{R}}^{t+1}_{I}]^{+}\|_{1}}=\frac{{\bm{R}}^{t+1}_{I}}{\|{\bm{R}}^{t+1}_{I}\|_{1}},$ (8)
where $i=P(I)$, and the second equality comes from the fact that ${\bm{R}}^{t+1}_{I}\geq\bm{0}$.

### v2

# Faster Game Solving via Asymmetry of Step Sizes

@@ Abstract
Counterfactual Regret Minimization (CFR) algorithms are widely used to compute a Nash equilibrium (NE) in two-player zero-sum imperfect-information extensive-form games (IIGs). Among them, Predictive CFR+ (PCFR+) is particularly powerful, achieving an exceptionally fast empirical convergence rate via the prediction in many games. However, the empirical convergence rate of PCFR+ would significantly degrade if the prediction is inaccurate, leading to unstable performance on certain IIGs. To enhance the robustness of PCFR+, we propose Asymmetric PCFR+ (APCFR+), which employs an adaptive asymmetry of step sizes between the updates of implicit and explicit accumulated counterfactual regrets to mitigate the impact of the prediction inaccuracy on convergence. We present a theoretical analysis demonstrating why APCFR+ can enhance the robustness. To the best of our knowledge, we are the first to propose the asymmetry of step sizes, a simple yet novel technique that effectively improves the robustness of PCFR+. Then, to reduce the difficulty of implementing APCFR+ caused by the adaptive asymmetry, we propose a simplified version of APCFR+ called Simple APCFR+ (SAPCFR+), which uses a fixed asymmetry of step sizes to enable only a single-line modification compared to original PCFR+. Experimental results on five standard IIG benchmarks and two heads-up no-limit Texas Hold’em (HUNL) Subagems show that (i) both APCFR+ and SAPCFR+ outperform PCFR+ in most of the tested games, (ii) SAPCFR+ achieves a comparable empirical convergence rate with APCFR+, and (iii) our approach can be generalized to improve other CFR algorithms, e.g., Discount CFR (DCFR).
Codehttps://github.com/menglinjian/AAAI-2026-APCFRPlus
@@ 1 Introduction
Figure 1: Comparison between PCFR+ and APCFR+, with differences highlighted in red. Note that the notation $t$ in $\alpha^{t}_{I}$ denotes iteration $t$, rather than an exponent.
Imperfect-information extensive-form games (IIGs) are foundational models to capture interactions among multiple agents in sequential settings with hidden information. IIGs are widely used to simulate real-world scenarios such as medical treatment (sandholm2015steering), security games (lisy2016counterfactual), cybersecurity (chen2017wireless), and recreational games (brown2018superhuman; brown2019superhuman). To address IIGs, a primary goal is to compute a Nash equilibrium (NE), where no player can unilaterally improve its payoff by deviating from the equilibrium.
As with much of the literature on solving IIGs, we focus on computing an NE in two-player zero-sum IIGs. The most widely used method for computing an NE in these IIGs is Counterfactual Regret Minimization (CFR)  (zinkevich2007regret; lanctot2009monte; tammelin2014solving; brown2019solving; farina2021faster; farina2019online; liu2021equivalence; liu2022power; meng2023efficient; farina2023regret; xu2022autocfr; xu2024minimizing; xu2024dynamic; zhang2024faster), as evidenced by their success in superhuman game AIs (bowling2015heads; moravvcik2017deepstack; brown2018superhuman; brown2019superhuman; perolat2022mastering). The key insight of CFR algorithms is to decompose the total regret over the game into a sum of counterfactual regrets associated within information sets (infosets) and employ a local regret minimizer to minimize counterfactual regrets within each infoset.
Many technologies have been proposed to improve the empirical convergence rate of CFR algorithms. For example, Counterfactual Regret Minimization+ (CFR+(tammelin2014solving) replaces the local regret minimizer—Regret Matching (RM) (hart2000simple; gordon2006no)—used in vanilla CFR with Regret Matching+ (RM+). CFR+ improves the empirical convergence rate by ensuring that the accumulated counterfactual regrets remain non-negative. Then, farina2021faster introduce Predictive CFR+ (PCFR+), an improved variant of CFR+.
PCFR+ significantly outperforms other CFR algorithms including CFR+ in many IIGs by using the prediction. Specifically, PCFR+ maintains two types of accumulated counterfactual regrets: the implicit and the explicit. As shown in Figure 1, at each iteration $t$, PCFR+ uses the prediction and the observed instantaneous counterfactual regret $\bm{r}^{t}_{I}$ to derive the new explicit accumulated counterfactual regret $\hat{\bm{R}}^{t}_{I}$ and the new implicit counterfactual regret $\bm{R}^{t+1}_{I}$, respectively. If the prediction aligns with the observed instantaneous counterfactual regret $\bm{r}^{t}_{I}$, the theoretical convergence rate of PCFR+ can be improved from $O(1/\sqrt{T})$ of CFR+ to $O(1/T)$ (farina2021faster). However, PCFR+ sets the instantaneous counterfactual regret $\bm{r}^{t-1}_{I}$ observed at iteration t1t-1 as the prediction at iteration $t$. This operation may cause inaccurate prediction on certain IIGs, which harms the empirical convergence rate of PCFR+. As noted by farina2021faster, PCFR+ underperforms other CFR algorithms in Leduc Poker (Game [O] in farina2021faster), yet significantly surpasses them in Battleship (3,2,3) (Game [R] in farina2021faster). This aligns with the results in Figure 2: the gap between predicted and observed instantaneous counterfactual regret decreases slowly in Leduc Poker but diminishes rapidly in Battleship (3,2,3), validating our hypothesis.
@@
Figure 2: Dynamics of inaccuracy in PCFR+ between predicted and observed instantaneous counterfactual regrets in Leduc Poker and Battleship (3,2,3). This inaccuracy is related to the theoretical convergence rate of PCFR+. The values on the y-axis are normalized to the range [0, 1], which is displayed on a logarithmic scale.
@@
To enhance the robustness of PCFR+, we propose a novel variant of PCFR+, termed Asymmetric PCFR+ (APCFR+). Similar to PCFR+, APCFR+ leverages the prediction to improve the convergence rate, but it mitigates the impact of the prediction inaccuracy on convergence. Specifically, as illustrated in Figure 1, APCFR+ utilizes an adaptive asymmetry mechanism for step sizes between implicit and explicit accumulated counterfactual regret updates, which dynamically reduces the step size when updating via the prediction. We prove that when the step size for updating the explicit accumulated counterfactual regret via the prediction at iteration $t$ is set to ${1}/{(1+\alpha^{t})}$ for APCFR+, where $\alpha^{t}\geq 0$ is a constant, the effect of the prediction inaccuracy on the convergence rate for APCFR+ is reduced by a factor of $1+\alpha^{t}$ compared to PCFR+. Therefore, APCFR+ mitigates the impact of the prediction inaccuracy on the convergence rate. Then, through the theoretical analysis of APCFR+, we propose an automatic learning mechanism for $\alpha^{t}$, eliminating the need for fine-tuning parameters across different games. To the best of our knowledge, we are the first to propose the asymmetry of step sizes updating implicit and explicit accumulated counterfactual regrets.
To simplify the implementation of APCFR+ caused by the automatic learning approach of $\alpha^{t}$, we introduce a simplified version of APCFR+, called Simple APCFR+ (SAPCFR+). Specifically, by analyzing the upper bounds of different terms within the theoretical guarantee of APCFR+ (detailed at the beginning of Section 4.2), SAPCFR+ sets $\alpha^{t}$ to $2$, ensuring SAPCFR+ requires only a single-line modification to the original PCFR+ code.
@@
We conduct extensive experimental evaluations of APCFR+ and SAPCFR+ across five standard IIG benchmarks as well as two heads-up no-limit Texas Hold’em (HUNL) Subgames generated by the top poker agent, Libratus (brown2018superhuman). The experiments demonstrate that APCFR+ and SAPCFR+ outperforms PCFR+ in nearly all tested games and achieve an empirical convergence rate comparable to that of PCFR+ in the remaining games. Moreover, we observe that SAPCFR+ achieves comparable empirical convergence rate with APCFR+. Finally, we can observe that our approach can be generalized to improve other CFR algorithms, e.g., Discount CFR (DCFR) (brown2019solving).
@@ 2 Related Work
We consider CFR algorithms (zinkevich2007regret; tammelin2014solving; brown2019solving; farina2021faster; farina2019online; liu2021equivalence; perolat2021poincare; liu2022power; meng2023efficient; farina2023regret; xu2022autocfr; xu2024minimizing; xu2024dynamic; zhang2024faster), the most widely used method for learning an NE in two-player zero-sum IIGs, as evidenced by their success in superhuman game AIs (bowling2015heads; moravvcik2017deepstack; brown2018superhuman; brown2019superhuman; perolat2022mastering).
The key insight of CFR algorithms is the decomposition of the regret over the game into the sum of counterfactual regrets associated with infosets. The vanilla CFR algorithm, introduced by zinkevich2007regret, employs RM (hart2000simple) as the local regret minimizer. To improve the empirical convergence rate of CFR, it is common to design more effective local regret minimizers, as the selection of the local regret minimizers has a significant impact on the overall performance of the CFR algorithm. Examples include RM+ (bowling2015heads), Discounted RM (DRM) (brown2019solving), and PRM+ (farina2021faster), which correspond to CFR+, DCFR, and PCFR+, respectively. PCFR+ can demonstrate an extremely faster empirical convergence rate than other CFR variants. However, as shown in its original paper, PCFR+ is outperformed by CFR+ and DCFR even on standard IIG benchmarks like Leduc Poker.
To improve the robustness of PCFR+, farina2023regret propose Stable PCFR+ and Smooth PCFR+. These algorithms improve the robustness by addressing the instability, i.e., rapid strategy fluctuations across iterations, via ensuring the lower bound of the 1-norm of accumulated counterfactual regrets exceeds a positive constant. However, these algorithms never outperform PCFR+ in terms of the empirical convergence rate even though they achieve a faster theoretical convergence rate than PCFR+, as demonstrated in our experiments. APCFR+ does not focus on addressing the instability, but instead aims to mitigate the impact of the prediction inaccuracy on the convergence to improve the robustness. In our experiments, APCFR+ consistently outperforms Stable PCFR+ and Smooth PCFR+ in all tested games.
@@ 3 Preliminaries
Imperfect-information Extensive-form games (IIGs). To model tree-form sequential decision-making problems with hidden information, a common used model is IIG (osborne2004introduction). An IIG can be formulated as $G=\{\mathcal{N},\mathcal{H},P,A,\mathcal{I},\{u_{i}\}\}$. Here, $\mathcal{N}=\{0,1\}$ is the set of players. “Nature” is also considered a player $c$ (representing chance) and chooses actions with a fixed known probability distribution. $\mathcal{H}$ is the set of all possible histories. For each history $h\in\mathcal{H}$, the function $P(h)$ represents the player acting at history $h$, and $A(h)$ denotes the actions available at history $h$. To account for private information, the histories for each player $i$ are partitioned into a collection $\mathcal{I}_{i}$, referred to as information sets (infosets). For any infoset $I\in\mathcal{I}_{i}$, histories $h,h^{\prime}\in I$ are indistinguishable to player $i$. The notation $\mathcal{I}$ denotes $\mathcal{I}=\{\mathcal{I}_{i}|i\in\mathcal{N}\}$. Thus, we have $P(I)=P(h)$, $A(I)=A(h),\forall h\in I$. The set of leaf nodes is denoted by $\mathcal{Z}$. For each leaf node $z$, there is a pair $(u_{0}(z),u_{1}(z))\in[-1,\ 1]$ which denotes the payoffs for the min player (player 0) and the max player (player 1) respectively. In two-player zero-sum IIGs, $u_{0}(z)=-u_{1}(z),\forall z\in\mathcal{Z}$.
@@
Behavioral strategy. This strategy $\sigma_{i}$ is defined on each infoset. For any infoset $I\in\mathcal{I}_{i}$, the probability for an action $a\in A(I)$ is denoted by $\sigma_{i}(I,a)$. We use $\sigma_{i}(I)=[\sigma_{i}(I,a)|a\in A(I)]\in\Delta^{|A(I)|}$ to denote the strategy at infoset $I$, where $\Delta^{|A(I)|}$ is a (|A(I)|1)(|A(I)|-1)-dimension simplex. If every player follows the strategy profile $\sigma=[\sigma_{0};\sigma_{1}]$ and reaches infoset $I$, the reaching probability is denoted by $\pi^{{\sigma}}(I)$. The probability contribution of player $i$ is $\pi_{i}^{\sigma}(I)$, while for players other than $i$, denoted as i-i, the contribution is $\pi_{-i}^{\sigma}(I)$. In IIGs, $u_{i}(\sigma)=u_{i}(\sigma_{i},\sigma_{-i})=\sum_{z\in\mathcal{Z}}u_{i}(z)\pi^{\sigma}(z)$.
@@
Nash equilibrium (NE). NE denotes a rational behavior where no player can benefit by unilaterally deviating from the equilibrium. For any player, her strategy is the best-response to the strategies of others. Formally, for any NE strategy profile $\sigma^{*}$ and $i\in\mathcal{N}$, it holds that $u_{i}(\sigma^{*}_{i},\sigma^{*}_{-i})\geq u_{i}(\sigma_{i},\sigma^{*}_{-i})$ for all $\sigma$. A widely used metric to measure the distance from the given strategy profile $\bm{x}$ to NE is the exploitability, which is defined as $\epsilon(\sigma)=\sum_{i\in\mathcal{N}}\max_{\sigma_{i}^{\prime}}(u_{i}(\sigma_{i}^{\prime},\sigma_{-i})-u_{i}(\sigma_{i},\sigma_{-i}))/|\mathcal{N}|$.
@@
Computing an NE via regret minimization algorithms. To compute an NE in IIGs, a common used method is regret minimization algorithms (rakhlin2013online; rakhlin2013optimization; hazan2016introduction; joulani2017modular). For any sequence of strategies $\sigma_{i}^{1},\cdots,\sigma_{i}^{T}$ of player $i$, player $i$’s regret is $R^{T}_{i}=\max_{\sigma_{i}}\sum^{T}_{t=1}u_{i}(\sigma_{i},\sigma^{t}_{-i})-\sum^{T}_{t=1}u_{i}(\sigma^{t}_{i},\sigma^{t}_{-i})$. Regret minimization algorithms are algorithms ensuring $R^{T}_{i}$ grows sublinearly. If each player follows a regret minimization algorithm, then their average strategy converges to the set of the NE in two-player zero-sum IIGs. Formally, assume the regret of each player $i$ is $R^{T}_{i}$, then it holds that
$\epsilon(\bar{\sigma})=\epsilon(\bar{\sigma}_{0},\bar{\sigma}_{1})\leq{\sum_{i\in\mathcal{N}}R^{T}_{i}}/{(|\mathcal{N}|T)},\$(2)
where $\bar{\sigma}_{i}(I)=\sum_{t=1}^{T}\pi^{\sigma^{t}}_{i}(I){\sigma^{t}_{i}}(I)/\sum_{t=1}^{T}\pi^{\sigma^{t}}_{i}(I)$.
@@
Counterfactual Regret Minimization (CFR) framework. This framework (zinkevich2007regret; farina2019online; liu2021equivalence) is designed to compute an NE of two-player zero-sum IIGs. Instead of directly minimizing the global regret $R^{T}_{i}$, it decomposes the regret to each infoset and independently minimizes the local regret within each infoset. Let $\sigma^{t}$ be the strategy profile at iteration $t$. This framework computes the counterfactual value at infoset $I$ for action $a$ as $v^{\sigma^{t}}(I,a)=\sum_{h\in I}\sum_{z\in\mathcal{Z}_{ha}}\pi^{\sigma^{t}}_{-i}(h)\pi^{\sigma^{t}}(ha,z)u_{i}(z),\$where $\pi^{\sigma^{t}}(ha,z)$ denotes the probability from $ha$ to $z$ if all players play according to $\sigma^{t}$ and $\mathcal{Z}_{ha}$ is the set of the leaf nodes that are reachable after choosing action $a$ at history $h$. For any infoset $I$, the counterfactual regret is $R^{T}(I)=\max_{a\in A(I)}\sum_{t=1}^{T}v^{\sigma^{t}}(I,a)-\sum_{t=1}^{T}\sum_{a^{\prime}\in A(I)}\sigma^{t}_{i}(I,a^{\prime})v^{\sigma^{t}}(I,a^{\prime}).$ The regret over the game $R^{T}_{i}=\max_{\sigma_{i}}\sum^{T}_{t=1}u_{i}(\sigma_{i},\sigma^{t}_{-i})-\sum^{T}_{t=1}u_{i}(\sigma^{t}_{i},\sigma^{t}_{-i})$ is less than the sum of the counterfactual regrets within infosets: $R^{T}_{i}\leq\sum_{I\in\mathcal{I}_{i}}R^{T}(I).$ So any regret minimization algorithms can be used as the local regret minimizer to minimize the regret $R^{T}(I)$ over each infoset to minimize the regret $R^{T}_{i}$.